# Operator Growth in Disordered Spin Chains: Indications for the Absence of Many-Body Localization

### v3

@@ III.1.1 No disorder
Note that the commutator has the form of a current density for $k$ odd and that of a charge density for $k$ even. From the formulas (III.1.1) and (III.1.1) we can immediately read off that there are $n(k)=\frac{k+1}{2}(k+1)\cdot 2=k+1+\frac{k}{2}(k+1)\cdot 2=(k+1)^{2}$ different terms generated at order $k$, both in the odd and in the even case. We confirm this result by calculating the commutator using symbolic manipulations up to high order, see Fig. 1.
Figure 1: XX model: Without disorder, the number of distinct terms $n(k)$ in $[H,\sigma^{z}_{0}]^{(k)}$ grows as $(k+1)^{2}$ while it grows according to Eq. (III.14) with disorder. The general bound (II.18), valid for any free-fermion model, is also shown.
Figure 2: In the XX model without disorder the norms grow as $s^{(1)}(k)=8^{k}$ and $s^{(2)}(k)=2^{k}\binom{2k}{k}$. Note that on this scale both results are very close to each other. The $2$-norm is only slightly smaller than the $1$-norm. Also shown is the general bound, Eq. (II.17), with $C=2$.
@@
$(C_{k})_{ij}=2^{k}{{k}\choose{i-1}}{{k}\choose{j-1}}$ (III.6)
with $11\leq i,j\leq k+$ 1. In the following, we concentrate on the case of $k$ odd which has a slightly simpler structure, see Eq. (III.1.1). In this case we find, in particular, that
$a_{kls}=\text{i}(-1)^{s}(C_{k})_{s+\frac{k+3}{2}+l,\frac{k+1}{2}-l}.$ (III.7)
From the result (III.7) we can now determine the total norms of the commutator and, furthermore, we can even determine which operators contribute how much to the total norm. From Eq. (III.1.1) and Eq. (III.7) we find that the set of the coefficients $\{|a_{kls}|\}$ is the same as $\{(C_{k})_{ij}\}^{2}$, i.e., each entry in $C_{k}$ appears two times. We therefore find
$s^{(1)}(k)\equiv||[H,\sigma^{z}_{0}]^{(k)}||_{1}=2\sum_{i=1}^{k+1}\sum_{j=1}^{\frac{k+1}{2}}(C_{k})_{ij}=8^{k}$ (III.8)
where we have used the explicit form of the coefficient matrix (III.6). Similarly, we find
$(s^{(2)}(k))^{2}\equiv||[H,\sigma^{z}_{0}]^{(k)}||^{2}_{2}=2\sum_{i=1}^{k+1}\sum_{j=1}^{\frac{k+1}{2}}(C_{k})_{ij}^{2}=2^{2k}{{2k}\choose{k}}^{2}.$ (III.9)
@@
$s^{(2)}(k)=2^{k}{{2k}\choose{k}}\approx\frac{8^{k}}{\sqrt{\pi k}}$ (III.10)
where we have used Stirling’s formula. This shows that the $2$-norm grows also exponentially in $k$ but with a correction which makes it always smaller than the $1$-norm as expected on general grounds, see Eq. (II). We can again check these results for finite $k$ by symbolic manipulations, see Fig. 2.
Finally, we can study how much terms which extend over a distance $l$ contribute to the total norm. We consider the case $k$ odd where the amplitude $a_{kls}$ for terms which have support on $22l+$ 2 sites is given by Eq. (III.7). We find again that the same entries appear multiple times and that all prefactors for a given $l$ are represented by the column $j=\frac{k+1}{2}-l$ of the coefficient matrix $(C_{k})_{ij}$. Summing over all initial sites we thus obtain
@@
$s^{(1)}_{l}(k)\approx\sqrt{\frac{8}{k\pi}}8^{k}\,.$ (III.12)
@@
Figure 3: For $k\gg l$ the norm of operators with support on $l$ sites is asymptotically described by Eq. (III.12) and is independent of $l$. The exact value (III.13) when $s_{l}^{(1)}(k)$ first becomes non-zero is also shown.
While we have been able to exactly calculate the $k$-th order commutator in the non-interacting, non-disordered case, see Eqs. (III.1.1, III.1.1, III.7), this is clearly no longer possible, even in principle, if random fields are included. On a positive note, this is also more than what we need. To distinguish between free-fermionic localized and non-localized phases and interacting localized and non-localized phases it is sufficient to consider the total commutator norm and the norm of the terms in the commutator which are supported on $l$ sites. This can be achieved more efficiently by developing graphs where the vertices represent the possible terms at order $k$ and the edges represent the contribution to the $1$-norm. Fig. 4 shows such a graph for the non-interacting model. The case without random magnetic fields discussed here corresponds to the case $D=0$ where only the terms in the left column are present.
@@
Figure 5: $s^{(1)}(k)$ and $s^{(2)}(k)$ for the Anderson model with $D=2$ and $D=10$. The results are averaged over $100$ samples. Note that the $2$-norm is only slightly smaller than the $1$-norm and the two norms are almost indistinguishable on this scale. The bound (III.15) is shown as well.
In this case, the graph is a $2$-regular graph with one node per level and $2$ edges per node. The constant number of nodes per level is directly reflected in the quadratic growth of the number of distinct terms $n(k)$.
To obtain $s^{(1)}(k)$, we have to sum over all weighted walks through the graph which start at the top and have length $k$ by multiplying the values of all traversed edges, irrespective of the vertex the walk ends on. If, on the other hand, we want to calculate only the contributions of terms which have support on $l$ sites, $s^{(1)}_{l}(k)$, then we have to sum only over those weighted walks with $k$ steps which end at a vertex with terms on $l$ sites. We discuss this approach to calculating or bounding the norm of the commutator in more detail in App. D. Fig. 4 also already shows how the graph has to be modified in the Anderson case which we discuss next.
III.1.2 Anderson Localization
Next, we consider the case of quenched disorder where the magnetic fields $h_{j}$ in the Hamiltonian (III.1) are drawn randomly from a box distribution, $h_{j}\in[-D,D]$. We note that in the one-dimensional case considered here any amount of disorder is a relevant perturbation. The model is localized for all $D>0$. Since $h_{j}$ is random, we cannot expect exact results for the norm. Instead, we will derive strict upper bounds.
@@ III.1.2 Anderson Localization
$n(k)=2k(k+1)+\left\{\begin{array}[]{cc}1&\mbox{even}\\ 1-k&\mbox{odd.}\end{array}\right.$ (III.14)
This result is compared to symbolic calculations in Fig. 1. We stress once more that Eq. (III.14) is independent of the type of disorder or the specific disorder configuration and is thus also valid in the Aubry-André case discussed in Sec. III.1.3. Note that while the precise number of terms has changed as compared to the non-disordered case, importantly the number of terms still only increases quadratically with $k$.
Figure 6: Left column: Norm ratio $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ for a fixed $k$ as a function of $l$ for the Anderson model with $D=2$ (top row) and $D=10$ (bottom row). Right column: $s^{(1)}_{l}(k)$ for a fixed $l$ as a function of $k$. The data are averaged over $100$ samples. Note that the value of $s_{l}^{(1)}(k)$ when it first becomes non-zero at $k=l-1$ is not affected by disorder, see Eq. (III.13).
@@
$s^{(1)}(k)\leq 8(8+8D)^{k-1}=8^{k}(D+1)^{k-1}.$ (III.15)
As expected, we find that the norm has an exponential bound. Note that this bound increases with increasing disorder $D$, see Fig. 5. Also note that for $D=0$ we obtain the exact result for the norm in the non-disordered case, Eq. (III.8). Apart from the initial node and the node on the top right, the graph with disorder is a $3$-regular graph. I.e., each node is connected to three edges. We note that in every $n$-regular graph the norm $s^{(1)}(k)$ can not grow faster than exponential. This remains true even if the number of nodes increases with $k$ as, for example, in a Cayley tree as long as the number of edges per node is constant. This follows immediately from the fact that $s^{(1)}(k)$ is the sum of the product of edge values along all possible paths of length $k$. That $n(k)$ grows only quadratically means that the graph in the free-fermion case is particularly simple and has a constant number of nodes at each level.
To see the difference between a free-fermionic model, where an initially localized operator spreads over the entire chain, and a system where the operator remains localized it is thus not sufficient to study the total operator norm. Instead, we have to investigate the spatial structure of the terms in the commutator at order $k$. Based on the graph in Fig. 4, we can derive a bound for the contribution of terms with support on $l$ sites to the norm at order $k$. We use again that $|h_{j}|\leq D$. The derivation of this bound is discussed in detail in App. D.2. Here, we will consider the simplest possible approach where we only consider those terms with support on $l$ sites which have the largest possible number of factors $h_{j}$. This approach is justified for $D\gg 1$ and, as we will argue, gives the same asymptotics in the limit $k\gg l\gg 1$ as the more sophisticated approach discussed in App. D.2. Let us first consider terms which have support on a single site, i.e., which are of the form $\sigma^{z}_{j}$. As is obvious from the graph in Fig. 4, such terms are only generated at even orders of $k$. We collect the largest number of factors $h_{j}$ if we go down one level, then travel $k-2$ times along the vertical direction, and then go back up to the top of the graph. The $1$-norm of such terms is thus bounded by
$s^{(1)}_{1}(k)\leq 8\times(8D)^{k-2}\times 4=2^{2k+1}(2D)^{k-2}\,.$ (III.16)
@@
$\frac{s_{l}(k)}{s_{1}(k)}\lesssim\binom{k-1}{l-2}(2D)^{3-l}\approx\frac{1}{\sqrt{2\pi l}}\left(\frac{ke}{l}\right)^{l}e^{-\frac{l^{2}}{2k}}(2D)^{3-l}\,.$ (III.18)
We note that while for a fixed $k$ we indeed find that the ratio (III.18) goes exponentially to zero for large $l$, there is a maximum in the ratio which shifts to larger $l$ with increasing $k$. This is not quite the behavior we expect. Instead, we expect that with increasing $k$ the ratio (III.18) converges to a function which is exponentially decreasing in $l$. Here it is important to note that (III.18) is only an upper bound. In particular, we have only considered those paths which have a maximum number of $h_{j}$ factors. There are, however, many additional paths and we expect that the prefactor in (III.18) will eventually converge with $k$ for a given $l$. This is consistent with what we observe from symbolic manipulations to finite orders of $k$, see Fig. 6. We note that in contrast to the non-disordered case where $s_{l}(k)$ grows exponentially and independent of $l$ for $k\gg l$, see Eq. (III.12) and Fig. 3, we find that with quenched disorder
$s_{l}(k)\sim\text{e}^{k}D^{-l}\,,$ (III.19)
@@
III.1.3 Aubry-André model
To further check that the exponential decay of $s_{l}(k)=||[H,\sigma^{z}_{0}]^{(k)}_{l}||$ with $l$ is indeed the signature of a localized system, we consider the Aubry-André model next. This is a model which is typically formulated as a fermionic model (III) where instead of quenched disorder we have a periodic potential
$h_{j}=\frac{\lambda}{2}\cos(2\pi\beta j+\phi)\,.$ (III.20)
Then, for $\beta$ Diophantine and almost any $\phi$—excluding a set of measure zero of resonant phases—the eigenstates of the Aubry-André model are localized for $\lambda>2$ [43, 44]. A common choice is to use the golden ratio $\beta=(1+\sqrt{5})/2$. The model thus provides an important test case where both a non-localized and a localized phase exist. Below the transition in the non-localized phase we expect that $s_{l}(k)$ is qualitatively described by Eq. (III.12), i.e., grows with $k$ independent of $l$ for $k\gg l$. In the localized phase, on the other hand, we expect that $s_{l}(k)$ is exponentially localized, see Eq. (III.19). Note that the graph for the Aubry-André model is the same as the one in the Anderson case shown in Fig. 4 with $D\to\lambda/2$ on the edges. With this replacement, also the bound (III.15) holds. We note that the considerations which let us to Eq. (III.19) in the Anderson case were valid only in the case of strong disorder. In this limit, they remain valid for the Aubry-André model. That the exponential decaying structure for the spatial support of the norm remains unchanged for all $D\neq 0$ in the Anderson case while there is a transition to the structure (III.12), which is independent of $l$ for $k\gg l$, for the Aubry-André model is beyond the strong disorder arguments we used in the previous section. All these results and expectations outlined above are fully consistent with the symbolic computations shown in Fig. 8. Note that the data in the left panels for the non-localized phase are shown on a linear scale while the results in the right panels for the localized phase are shown on a logarithmic scale.
@@ III.2 Interacting case
With the results we have established in the non-interacting non-localized and localized cases, we can now look at the interacting isotropic Heisenberg (XXX) model, i.e., the Hamiltonian (III.1) with $\Delta=1$. A fundamental general difference in interacting as compared to non-interacting models is that one-body operators do not remain one-body operators when commuted with the Hamiltonian. This leads to a qualitatively different phenomenology in the operator growth with exponentially many more terms and paths.
@@
We also note that the Heisenberg model without disorder is a so-called integrable model which has, what is commonly called, an infinite set of local conserved charges $O$ with $[H,O]=0$ in the thermodynamic limit. It is crucial to note though that ‘local’ is meant here in an entirely different way. These charges are in no way localized but rather can be written as $O=\sum_{j}o_{j}$ where $o_{j}$ is a local density acting on a finite number of lattice sites. I.e., the charge densities are local but the entire conserved operator $O$ is not and rather acts on the entire lattice. The trivial examples are the Hamiltonian itself, $H=\sum_{I}h_{I}$ and the magnetization $M=\sum_{I}\sigma^{z}_{I}$. What is special about integrable models is that an infinite set of operators of this type exist which allows to uniquely characterize all eigenstates by the corresponding quantum numbers. The eigenspectrum of the model can thus be calculated exactly by using the Bethe ansatz [45, 46]. For the question of local operator growth, integrability is not relevant and we do not expect to see any difference to a non-integrable model. Even if we break integrability, the magnetization $M$ is still conserved as long as $U(1)$ symmetry is preserved. I.e., we are in any case studying the operator growth of a local operator $\sigma^{z}_{j}$ which belongs to a globally conserved charge. Differences between integrable and non-integrable model are, of course, expected if we study the Euclidean time dynamics of operators of the type $O=\sum_{j}o_{j}$ [31]. However, for the Euclidean time dynamics of local operators in a generic system, the only thing which matters is whether or not this system has local conserved charges, i.e., whether it is localized or not. We also note that integrability is immediately destroyed if we introduce even an infinitesimal amount of disorder.
@@
We will first briefly consider the case without disorder before turning to the most interesting case, the XXX model with random magnetic fields $h_{j}\in[-D,D]$ which has been suggested to contain a localized phase for sufficiently strong disorder.
III.2.1 No disorder

### v4

@@ V.1.1 No disorder
Note that the commutator has the form of a current density for $k$ odd and that of a charge density for $k$ even. From the formulas (V.1.1) and (V.1.1) we can immediately read off that there are $n(k)=\frac{k+1}{2}(k+1)\cdot 2=k+1+\frac{k}{2}(k+1)\cdot 2=(k+1)^{2}$ different terms generated at order $k$, both in the odd and in the even case. We confirm this result by calculating the commutator using symbolic manipulations up to high order, see Fig. 4.
Figure 4: XX model: Without disorder, the number of distinct terms $n(k)$ in $[H,\sigma^{z}_{0}]^{(k)}$ grows as $(k+1)^{2}$ while it grows according to Eq. (V.14) with disorder. The general bound (III.18), valid for any free-fermion model, is also shown.
Figure 5: In the XX model without disorder the norms grow as $s^{(1)}(k)=8^{k}$ and $s^{(2)}(k)=2^{k}\binom{2k}{k}$. Note that on this scale both results are very close to each other. The $2$-norm is only slightly smaller than the $1$-norm. Also shown is the general bound, Eq. (III.17), with $C=2$.
@@
$(C_{k})_{ij}=2^{k}{{k}\choose{i-1}}{{k}\choose{j-1}}$ (V.6)
with $11\leq i,j\leq k+$ 1. In the following, we concentrate on the case of $k$ odd which has a slightly simpler structure, see Eq. (V.1.1). In this case we find, in particular, that
$a_{kls}=\text{i}(-1)^{s}(C_{k})_{s+\frac{k+3}{2}+l,\frac{k+1}{2}-l}.$ (V.7)
From the result (V.7) we can now determine the total norms of the commutator and, furthermore, we can even determine which operators contribute how much to the total norm. From Eq. (V.1.1) and Eq. (V.7) we find that the set of the coefficients $\{|a_{kls}|\}$ is the same as $\{(C_{k})_{ij}\}^{2}$, i.e., each entry in $C_{k}$ appears two times. We therefore find
$s^{(1)}(k)\equiv||[H,\sigma^{z}_{0}]^{(k)}||_{1}=2\sum_{i=1}^{k+1}\sum_{j=1}^{\frac{k+1}{2}}(C_{k})_{ij}=8^{k}$ (V.8)
where we have used the explicit form of the coefficient matrix (V.6). Similarly, we find
$(s^{(2)}(k))^{2}\equiv||[H,\sigma^{z}_{0}]^{(k)}||^{2}_{2}=2\sum_{i=1}^{k+1}\sum_{j=1}^{\frac{k+1}{2}}(C_{k})_{ij}^{2}=2^{2k}{{2k}\choose{k}}^{2}.$ (V.9)
@@
$s^{(2)}(k)=2^{k}{{2k}\choose{k}}\approx\frac{8^{k}}{\sqrt{\pi k}}$ (V.10)
where we have used Stirling’s formula. This shows that the $2$-norm grows also exponentially in $k$ but with a correction which makes it always smaller than the $1$-norm as expected on general grounds, see Eq. (III). We can again check these results for finite $k$ by symbolic manipulations, see Fig. 5.
Finally, we can study how much terms which extend over a distance $l$ contribute to the total norm. We consider the case $k$ odd where the amplitude $a_{kls}$ for terms which have support on $22l+$ 2 sites is given by Eq. (V.7). We find again that the same entries appear multiple times and that all prefactors for a given $l$ are represented by the column $j=\frac{k+1}{2}-l$ of the coefficient matrix $(C_{k})_{ij}$. Summing over all initial sites we thus obtain
@@
$s^{(1)}_{l}(k)\approx\sqrt{\frac{8}{k\pi}}8^{k}\,.$ (V.12)
@@
Figure 6: For $k\gg l$ the norm of operators with support on $l$ sites is asymptotically described by Eq. (V.12) and is independent of $l$. The exact value (V.13) when $s_{l}^{(1)}(k)$ first becomes non-zero is also shown.
While we have been able to exactly calculate the $k$-th order commutator in the non-interacting, non-disordered case, see Eqs. (V.1.1, V.1.1, V.7), this is clearly no longer possible, even in principle, if random fields are included. On a positive note, this is also more than what we need. To distinguish between free-fermionic localized and non-localized phases and interacting localized and non-localized phases it is sufficient to consider the total commutator norm and the norm of the terms in the commutator which are supported on $l$ sites. This can be achieved more efficiently by developing graphs where the vertices represent the possible terms at order $k$ and the edges represent the contribution to the $1$-norm. Fig. 7 shows such a graph for the non-interacting model. The case without random magnetic fields discussed so far corresponds to the case $D=0$ where only the terms in the left column are present.
@@
Figure 8: $s^{(1)}(k)$ and $s^{(2)}(k)$ for the Anderson model with $D=2$ and $D=10$. The results are averaged over $100$ samples. Note that the $2$-norm is only slightly smaller than the $1$-norm and the two norms are almost indistinguishable on this scale. The bound (V.15) is shown as well.
In this case, the graph is a $2$-regular graph with one node per level and two edges per node. The constant number of nodes per level is directly reflected in the quadratic growth of the number of distinct terms $n(k)$.
To obtain $s^{(1)}(k)$, we have to sum over all weighted walks through the graph which start at the top and have length $k$ by multiplying the values of all traversed edges, irrespective of the vertex the walk ends on. If, on the other hand, we want to calculate only the contributions of terms which have support on $l$ sites, $s^{(1)}_{l}(k)$, then we have to sum only over those weighted walks with $k$ steps which end at a vertex with terms on $l$ sites. We discuss this approach to calculating or bounding the norm of the commutator in more detail in the Suppl. Mat. Fig. 7 also already shows how the graph has to be modified in the Anderson case which we discuss next.
V.1.2 Anderson Localization
Next, we consider the case of quenched disorder where the magnetic fields $h_{j}$ in the Hamiltonian (V.1) are drawn randomly from a box distribution, $h_{j}\in[-D,D]$. We note that in the one-dimensional case considered here any amount of disorder is a relevant perturbation. The model is localized for all $D>0$. Since $h_{j}$ is random, we cannot expect exact results for the norm. Instead, we will derive strict upper bounds.
@@ V.1.2 Anderson Localization
$n(k)=2k(k+1)+\left\{\begin{array}[]{cc}1&\mbox{even}\\ 1-k&\mbox{odd.}\end{array}\right.$ (V.14)
This result is compared to symbolic calculations in Fig. 4. We stress once more that Eq. (V.14) is independent of the type of disorder or the specific disorder configuration and is thus also valid in the Aubry-André case discussed in Sec. V.1.3. Note that while the precise number of terms has changed as compared to the non-disordered case, importantly the number of terms still only increases quadratically with $k$.
Figure 9: Left column: Norm ratio $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ for a fixed $k$ as a function of $l$ for the Anderson model with $D=2$ (top row) and $D=10$ (bottom row). Right column: $s^{(1)}_{l}(k)$ for a fixed $l$ as a function of $k$. The data are averaged over $100$ samples. Note that the value of $s_{l}^{(1)}(k)$ when it first becomes non-zero at $k=l-1$ is not affected by disorder, see Eq. (V.13).
@@
$s^{(1)}(k)\leq 8(8+8D)^{k-1}=8^{k}(D+1)^{k-1}.$ (V.15)
As expected, we find that the norm has an exponential bound. Note that this bound increases with increasing disorder $D$, see Fig. 8. Also note that for $D=0$ we obtain the exact result for the norm in the non-disordered case, Eq. (V.8). Apart from the initial node and the node on the top right, the graph with disorder is a $3$-regular graph. I.e., each node is connected to three edges. We note that in every $n$-regular graph the norm $s^{(1)}(k)$ can not grow faster than exponential. This remains true even if the number of nodes increases with $k$ as, for example, in a Cayley tree as long as the number of edges per node is constant. This follows immediately from the fact that $s^{(1)}(k)$ is the sum of the product of edge values along all possible paths of length $k$. That $n(k)$ grows only quadratically means that the graph in the free-fermion case is particularly simple and has a constant number of nodes at each level.
To see the difference between a free-fermionic model, where an initially localized operator spreads over the entire chain, and a system where the operator remains localized it is thus not sufficient to study the total operator norm. Instead, we have to investigate the spatial structure of the terms in the commutator at order $k$. Based on the graph in Fig. 7, we can derive a bound for the contribution of terms with support on $l$ sites to the norm at order $k$. We use again that $|h_{j}|\leq D$. The derivation of this bound is discussed in detail in the Suppl. Mat. Here, we will consider the simplest possible approach where we only take into account those terms with support on $l$ sites which have the largest possible number of factors $h_{j}$. This approach is justified for $D\gg 1$ and, as we will argue, gives the same asymptotics in the limit $k\gg l\gg 1$ as the more sophisticated approach discussed in the Suppl. Mat. Let us first consider terms which have support on a single site, i.e., which are of the form $\sigma^{z}_{j}$. As is obvious from the graph in Fig. 7, such terms are only generated at even orders of $k$. We collect the largest number of factors $h_{j}$ if we go down one level, then travel $k-2$ times along the vertical direction, and then go back up to the top of the graph. The $1$-norm of such terms is thus bounded by
$s^{(1)}_{1}(k)\leq 8\times(8D)^{k-2}\times 4=2^{2k+1}(2D)^{k-2}\,.$ (V.16)
@@
$\frac{s_{l}(k)}{s_{1}(k)}\lesssim\binom{k-1}{l-2}(2D)^{3-l}\approx\frac{1}{\sqrt{2\pi l}}\left(\frac{ke}{l}\right)^{l}e^{-\frac{l^{2}}{2k}}(2D)^{3-l}\,.$ (V.18)
We note that while for a fixed $k$ we indeed find that the ratio (V.18) goes exponentially to zero for large $l$, there is a maximum in the ratio which shifts to larger $l$ with increasing $k$. This is not quite the behavior we expect. Instead, we expect that with increasing $k$ the ratio (V.18) converges to a function which is exponentially decreasing in $l$. Here it is important to note that (V.18) is only an upper bound. In particular, we have only considered those paths which have a maximum number of $h_{j}$ factors. There are, however, many additional paths and we expect that the prefactor in (V.18) will eventually converge with $k$ for a given $l$. This is consistent with what we observe from symbolic manipulations to finite orders of $k$, see Fig. 9. We note that in contrast to the non-disordered case where $s_{l}(k)$ grows exponentially and independent of $l$ for $k\gg l$, see Eq. (V.12) and Fig. 6, we find that with quenched disorder
$s_{l}(k)\sim\text{e}^{k}D^{-l}\,,$ (V.19)
@@
V.1.3 Aubry-André model
To further check that the exponential decay of $s_{l}(k)=||[H,\sigma^{z}_{0}]^{(k)}_{l}||$ with $l$ is indeed the signature of a localized system, we consider the Aubry-André model next. This is a model which is typically formulated as a fermionic model (V) where instead of quenched disorder we have a periodic potential
$h_{j}=\frac{\lambda}{2}\cos(2\pi\beta j+\phi)\,.$ (V.20)
Then, for $\beta$ Diophantine and almost any $\phi$—excluding a set of measure zero of resonant phases—the eigenstates of the Aubry-André model are localized for $\lambda>2$ [44, 45]. A common choice is to use the golden ratio $\beta=(1+\sqrt{5})/2$. The model thus provides an important test case where both a non-localized and a localized phase exist. Below the transition in the non-localized phase we expect that $s_{l}(k)$ is qualitatively described by Eq. (V.12), i.e., grows with $k$ independent of $l$ for $k\gg l$. In the localized phase, on the other hand, we expect that $s_{l}(k)$ is exponentially localized, see Eq. (V.19). Note that the graph for the Aubry-André model is the same as the one in the Anderson case shown in Fig. 7 with $D\to\lambda/2$ on the edges. With this replacement, also the bound (V.15) holds. We note that the considerations which let us to Eq. (V.19) in the Anderson case were valid only in the case of strong disorder. In this limit, they remain valid for the Aubry-André model. That the exponential decaying structure for the spatial support of the norm remains unchanged for all $D\neq 0$ in the Anderson case while there is a transition to the structure (V.12), which is independent of $l$ for $k\gg l$, for the Aubry-André model is beyond the strong disorder arguments we used in the previous section. All these results and expectations outlined above are fully consistent with the symbolic computations shown in Fig. 11. Note that the data in the left panels for the non-localized phase are shown on a linear scale while the results in the right panels for the localized phase are shown on a logarithmic scale.
@@ V.2 Interacting case
After having checked the analytical results derived in Sec. III and Sec. IV for the non-interacting non-localized and localized cases, we can now look at the interacting isotropic Heisenberg (XXX) model, i.e., the Hamiltonian (V.1) with $\Delta=1$. A fundamental general difference in interacting as compared to non-interacting models is that one-body operators do not remain one-body operators when commuted with the Hamiltonian. This leads to a qualitatively different phenomenology in the operator growth with exponentially many more terms and paths.
@@
We also note that the Heisenberg model without disorder is a so-called integrable model which has, what is commonly called, an infinite set of local conserved charges $O$ with $[H,O]=0$ in the thermodynamic limit. It is crucial to note though that ‘local’ is meant here in an entirely different way. These charges are in no way localized but rather can be written as $O=\sum_{j}o_{j}$ where $o_{j}$ is a local density acting on a finite number of lattice sites. I.e., the charge densities are local but the entire conserved operator $O$ is not and rather acts on the entire lattice. The trivial examples are the Hamiltonian itself, $H=\sum_{I}h_{I}$ and the magnetization $M=\sum_{I}\sigma^{z}_{I}$. What is special about integrable models is that an infinite set of operators of this type exist which allows to uniquely characterize all eigenstates by the corresponding quantum numbers. The eigenspectrum of the model can thus be calculated exactly by using the Bethe ansatz [46, 47]. For the question of local operator growth, integrability is not relevant and we do not expect to see any difference to a non-integrable model. Even if we break integrability, the magnetization $M$ is still conserved as long as $U(1)$ symmetry is preserved. I.e., we are in any case studying the operator growth of a local operator $\sigma^{z}_{j}$ which belongs to a globally conserved charge. Differences between integrable and non-integrable models are, of course, expected if we study the Euclidean time dynamics of operators of the type $O=\sum_{j}o_{j}$ [31]. However, for the Euclidean time dynamics of local operators in a generic system, the only thing which matters is whether or not this system has local conserved charges, i.e., whether it is localized or not. We also note that integrability is immediately destroyed if we introduce even an infinitesimal amount of disorder.
@@
We will first briefly consider the case without disorder before turning to the most interesting case, the XXX model with random magnetic fields $h_{j}\in[-D,D]$ which has been suggested to have a localized phase for sufficiently strong disorder.
V.2.1 No disorder